OpenStudy (anonymous):

(5ab - 10b 2 + 15bc) ÷ 5b

OpenStudy (tyteen4a03):

\(\frac{(5ab - 10b^2 + 15bc)}{5b}\) Notice how the variable b exists across every variable.

OpenStudy (aravindg):

@lilpeabody do you observe any factor that we can take out from numerator?

Parth (parthkohli):

Also look at some cool stuff like this:\[\dfrac{\rm blah + bleh + bloh}{\rm burn} = \rm \dfrac{blah}{burn} + \dfrac{bleh}{burn} + \dfrac{bloh}{burn}\]Look Charlie, everything's burnt. T_T